OpenStudy (anonymous):

the angle of elevation from a point on the ground to the top of the tower id 36 degrees 6 minutes. the angle of elevation from a point 192 feet farther back is 27 degrees 32 minutes. find the height of the tower. round answer to nearest hundredth.

OpenStudy (xapproachesinfinity):



OpenStudy (xapproachesinfinity):

this the diagram see what can we get from this

OpenStudy (anonymous):

im looking for the height of the tower

OpenStudy (xapproachesinfinity):

i'm lagging ...

OpenStudy (xapproachesinfinity):

that the height of the tower i called h

OpenStudy (anonymous):

oh I see

OpenStudy (anonymous):

x is 140?

OpenStudy (anonymous):

I used tangent ratio with tan 36.1=x/192

OpenStudy (xapproachesinfinity):

hmm how did you use it

OpenStudy (anonymous):

tan36.1 =x/192 x=192tan(36.1) x=192(0.729)

OpenStudy (xapproachesinfinity):

we have two equation involving tan \[\Large \begin {cases} \tan(27^{\circ}32')=\frac{h}{x}\\ \tan (36^{\circ}6')=\frac{h}{x+192}\end{cases}\]

OpenStudy (xapproachesinfinity):

of course you need to convert those angle measure to only degree first

OpenStudy (xapproachesinfinity):

just remember that 60 min = 1 rotation = 360 degrees

OpenStudy (xapproachesinfinity):

or just 30 min = 180 degrees

OpenStudy (anonymous):

idfk?

OpenStudy (xapproachesinfinity):

just do \[27 + 32/60\]

OpenStudy (xapproachesinfinity):

and the same for the other one

OpenStudy (xapproachesinfinity):

sorry the net is bad, i'm lagging

OpenStudy (xapproachesinfinity):

36 + 6/60 the result will be in degrees only

OpenStudy (xapproachesinfinity):

after that solve using the two equations

OpenStudy (anonymous):

I already know 36 degrees 6 minutes is 36.1 and 27 degrees is 27.53 what next

OpenStudy (xapproachesinfinity):

next is to evaluate tan of those angles

OpenStudy (xapproachesinfinity):

\[0.52=\frac{h}{x}\] \[0.73=\frac{h}{x+192}\]

OpenStudy (xapproachesinfinity):

solve for x in the first equation then replace what you found into the second equation this is basic system of two equations

OpenStudy (xapproachesinfinity):

\[x=\frac{h}{0.52} \Longrightarrow 0.73=\frac{h}{\frac{h}{0.52}+192}\]